OpenStudy (anonymous):

Find all solutions if 0° ≤ θ < 360°. When necessary, round your answers to the nearest tenth of a degree. (Enter your answers as a comma-separated list.) cos θ + sin θ = 1

OpenStudy (paxpolaris):

\[\sin \theta+\cos \theta= \sqrt2\cos \left( \theta-\frac \pi4 \right)\]

OpenStudy (paxpolaris):

at 0°: sin theta = 0, cos theta =1 at 90°: sin theta = 1, cos theta =0